OpenStudy (anonymous):

WILL GIVE MEDALS find all zeros of the equation -3x^4+27x^2+1200=0

OpenStudy (anonymous):

@hartnn pls help

OpenStudy (anonymous):

\[-3(x^4 - 9x^2 - 400) = 0\] \[-3(x^2 - 25)(x^2 + 16) = 0\] Set each factor equal to zero. \[x^2 - 25 = 0\] \[x = 5, -5\] \[x^2 + 16 = 0\] \[x = 4i, -4i\] Leave -3 out because when you set =3 equal to zero: -3 = 0 they are NOT equal to zero!!!!! But when you simplify, you should write it out as: \[-3(x^2 + 16)(x-5)(x+5) = 0\]

hartnn (hartnn):

got that ? still need any help ?

OpenStudy (anonymous):

nope im good, thanks skiller!

OpenStudy (anonymous):

Glad I could assist you :)